Solve the following equation. V5x + 21 - 3 = x (Type an integer or a simplified fraction.
Answer #1

Given that, 5x+31 - 3 = X i.e. 5x+ 21 X + 3 e. (152+27) = (x+3) je, 57+21 = x²+6x+9 ice, +6% +9-52-91 = 0 lie. x²+x x +X-12 =​​​​​​.

.

So the solution is x = 3